Given the demand function
Qd=650-5p-p2(p square) where p= 10
Determine the price elasticity of demand.
1
Expert's answer
2018-04-12T09:37:11-0400
e = (P/Q)*(dQ/dP) at p=10
dQ/dP(p=10) = -5 - 2p = -25
e = -10*25/650 = -0.38
So, this is relatively inelastic demand